# Underwater Image Quality Assessment: A Perceptual Framework Guided by Physical Imaging

### v1

@@ Abstract
In this paper, we propose a physically imaging-guided framework for underwater image quality assessment (UIQA), called PIGUIQA. First, we formulate UIQA as a comprehensive problem that considers the combined effects of direct transmission attenuation and backwards scattering on image perception. On this basis, we incorporate advanced physics-based underwater imaging estimation into our method and define distortion metrics that measure the impact of direct transmission attenuation and backwards scattering on image quality. Second, acknowledging the significant content differences across various regions of an image and the varying perceptual sensitivity to distortions in these regions, we design a local perceptual module on the basis of the neighborhood attention mechanism. This module effectively captures subtle features in images, thereby enhancing the adaptive perception of distortions on the basis of local information. Finally, by employing a global perceptual module to further integrate the original image content with underwater image distortion information, the proposed model can accurately predict the image quality score. Comprehensive experiments demonstrate that PIGUIQA achieves state-of-the-art performance in underwater image quality prediction and exhibits strong generalizability. The code for PIGUIQA is available on https:// anonymous.4open.science/r/PIGUIQA-A465/.
@@ 1 Introduction
The underwater world is rich in resources, and underwater images provide a crucial medium for exploring environments such as oceans and lakes by accurately and intuitively capturing underwater information. However, the complexities of the underwater imaging environment often result in suboptimal image quality [14, 47]. Therefore, underwater image quality assessment (UIQA) is fundamental for underwater image processing. UIQA methods help determine whether underwater images meet quality requirements for various applications and assess the visual quality of enhanced underwater images.
According to the accessibility of ideal references, image quality assessment (IQA) methods can be broadly categorized into full-reference (FR) and no-reference (NR) approaches. For underwater images, obtaining perfect undistorted references is often impractical; thus, typical FR-IQA approaches, such as the structural similarity index measure (SSIM) [36] and feature similarity indexing method (FSIM) [43], are not applicable. Although NR-IQA approaches [26, 7, 29] have been developed for many years, most of them are designed to evaluate generic images. Many NR-IQA methods [21, 28, 19] are based on statistical properties, such as natural scene statistics (NSS), which are typically observed under good lighting conditions and in clear visual environments. However, the unique characteristics of underwater environments result in light propagation and scattering rules that are significantly different from those in terrestrial natural scenes. As a result, the statistical properties of NSSs often fall short in accurately describing and assessing the quality of underwater images.
@@
Another primary challenge is that underwater images differ markedly from terrestrial images because of their specific imaging environments and lighting conditions. As a result, traditional NR-IQA methods that perform well in terrestrial image processing cannot be directly applied to underwater images. Specifically, as shown in Fig. 1, the underwater environment is complex and variable, with light propagation affected by scattering and absorption. Different wavelengths of light attenuate at different rates, with longer wavelengths diminishing more rapidly in water. Consequently, shorter wavelengths such as green and blue light can travel further underwater, causing underwater images to exhibit a blue–green hue and suffering from significant color distortion. The scattering of light in water can be categorized into forward scattering and backwards scattering on the basis of the angle. Forward scattering occurs when light reflected from the target object deviates from its original path before it reaches the sensor, resulting in blurred underwater images. Backwards scattering, on the other hand, involves a large amount of stray light entering the sensor, leading to visual issues such as loss of detail, low overall contrast, and increased background noise in underwater images. The need for the development of specialized UIQA methods that can effectively address the unique challenges posed by underwater imaging is urgent.
@@
To address the aforementioned issues, combining prior knowledge of underwater imaging with the excellent feature extraction capabilities of deep learning presents an effective solution. On this basis, we propose a physically imaging-guided framework for UIQA (PIGUIQA) by integrating the physical model of underwater light propagation with deep learning techniques. Specifically, the framework considers distortions caused by backwards scattering and direct transmission processes. These factors play crucial roles in underwater imaging, influencing image clarity, contrast, and color representation. Within the PIGUIQA framework, we leverage the powerful feature extraction capabilities of deep learning models to capture both local details and global perceptual features in images. Capturing local details allows the model to perceive distortions in critical regions of underwater images, such as object edges and texture information, whereas global perceptual features help the model understand the overall structure and context of the entire image. The main contributions of this work can be summarized as follows:
@@
  • We formulate UIQA as a comprehensive problem that accounts for the combined effects of direct transmission attenuation and backwards scattering on image perception. We integrate advanced physics-based underwater imaging estimations into our framework. From the perspective of matrix theory, we define distortion metrics to measure these impacts.
@@
  • We design a neighborhood attention (NA)-based local perceptual module to capture subtle features in images, thereby enhancing the adaptive perception of distortions via local information. Additionally, a global perceptual module is employed to integrate the original image content with underwater image distortion information, assisting the model in understanding the overall structure and context of the entire image.
@@ 3.1 Problem formulation
For underwater images, we specifically analyse image quality by considering the causes of distortion inherent to underwater imaging. In real underwater scenarios, the camera is often positioned close to the underwater scene. Consequently, the impact of forward scattering on image quality can be neglected, resulting in only direct transmission attenuation and backwards scattering components to be considered. This relationship is represented as:
@@
where $Q$ is the ground-truth quality of an underwater image and where $d_{1}$ and $d_{2}$ represent the distance maps measuring transmission attenuation and backwards scattering distortion between $T$ and $E$ and between $B$ and $O$, respectively. $f_{1}$ and $f_{2}$ are local distortion-aware functions to be learned, and $f_{3}$ is a global perceptual function to be learned. This formulation captures the essential aspects of UIQA by integrating a physical model into the evaluation process.
@@ 3.2 Framework
3.3 Underwater Imaging Estimation
A commonly used precise revised underwater image formation model [1] is as follows:
@@ 3.3 Underwater Imaging Estimation
where $c\in\{R,G,B\}$ is the color channel; $I_{per}$ and $I_{dis}$ are distortion-free and distorted images, respectively; $l$ is the object‒camera distance; $B_{\infty}$ is the background ambient light image; and $\beta_{D}$ and $\beta_{B}$ are the direct attenuation and backwards scattering coefficients, respectively.
In this paper, we utilize the SyreaNet model proposed by Wen et al. [38] for underwater imaging estimation, which is represented mathematically as follows:
$\hat{T},\hat{B},\hat{I}_{per}=SyreaNet(I_{dis})$ (5)
This method synthesizes more formal underwater images under various conditions via the imaging model described in Eq. (4). Subsequently, it trains on the synthesized underwater image data to estimate the transmission attenuation matrix $T$ and background clutter matrix $B$, as defined in Eq. (2). Compared with traditional manual estimation of underwater imaging parameters, which often lacks accuracy, SyreaNet integrates physical modelling with the interpretability of imaging models and the feature extraction capabilities of neural networks. As illustrated in Fig. 3, it can automatically learn and extract features, adaptively incorporating prior knowledge on the basis of different underwater environments and imaging conditions. This results in improved adaptability and robustness in underwater imaging estimation.
Figure 3: Illustration of underwater imaging estimation by SyreaNet [38]. (a): Input distorted underwater images $I_{dis}$. (b): Estimated transmission attenuation map $\hat{T}$. (c): Estimated background clutter map $\hat{B}$. (d): Restored underwater images $\hat{I}_{per}$. In the transmission attenuation map, darker areas signify more severe scattering and greater distortion, whereas in the background clutter map, brighter areas indicate a greater degree of background noise.
3.4 Imaging Distortion Measurements
Considering the varying depths and optical properties of different regions and objects in underwater images, we decompose the image transmission attenuation map $T$ and the background clutter map $B$ into patches of size $N\times N$. This strategy facilitates the model’s learning of local features, thereby enhancing the perception of details in images with intricate content. Consequently, we address the design of imaging distortion metrics for each patch.
@@ 3.4 Imaging Distortion Measurements
The background clutter map can be regarded as an additional noise overlay on the original image, arising from background textures, variations in lighting, or the presence of nontarget objects. According to the theory of human visual sensitivity, the HVS can only distinctly recognize distortion when the intensity of the noise surpasses a specific perceptual threshold. Thus, we propose a backwards scattering distortion measurement for the $k$-th patch of $B$:
@@ 4.1 Experimental Settings
WaDIQaM [3] $\mathbf{0.8565}$ 0.8408 $\mathbf{0.6617}$ $\mathbf{10.8477}$ $\mathbf{0.5099}$ $\mathbf{0.4876}$ $\mathbf{0.3414}$ $\mathbf{1.8250}$
B-FEN [39] 0.8512 $\mathbf{0.8441}$ 0.6595 10.8642 0.4841 0.4766 0.3261 1.8564
@@
Implementation details. The proposed FIGUIQA network is trained ***The implementation code is programmed in PyTorch and is publicly available at https://anonymous.4open.science/r/PIGUIQA-A465/. The experiments are run on a PC with a Linux operating system and are configured with an NVIDIA GeForce RTX 4090 GPU with 24 GB of memory. using the default settings of the Adam optimizer. The learning rate is set to $10^{-4}$. To ensure the stability of the final model, a total of 1000 epochs are conducted during training. To prevent overfitting, data augmentation techniques such as random horizontal and vertical flipping, as well as rotation within the range of $[-15,15]$ degrees, are employed. The size of patches is set to $16\times 16$, which is a common setting in computer vision.
@@ 4.2 Performance Comparison
We compare the proposed method against several commonly used no-reference image quality assessment (NR-IQA) techniques and state-of-the-art user-in-the-loop image quality assessment (UIQA) methods, including BRISQUE [26], BLIINDSII [32], NFERM [7], NIQE [27], IL-NIQE [44], SNP-NIQE [22], BIQME [8], FRIQUEE [5], PIQE [29], NRSL [19], RISE [18], DIIVINE [28], NPQI [21], dipIQ [24], HyperIQA [33], WaDIQaM, UNIQUE [46], DBCNN [45], B-FEN [39], UCIQE [40], UIQM [30], CCF [35], FDUM [42], NUIQ [15], MCOLE [16], Twice-Mix [4], and UIQI [23]. To ensure the reliability and stability of the proposed NAFRAD method, we repeat the training and testing process 10 times, using the average of the test results as the final metric.
@@ 4.2.1 Single-Dataset Scenarios
In the experimental scenario involving single datasets, we randomly selected 80% of the data from each dataset for training and used the remaining 20% for evaluation. The experimental results are presented in Tables 1, 2, and LABEL:tab:performance3. Our proposed method achieves the highest performance on the SAUD2.0 dataset, significantly surpassing the second-place method, MCOLE [16], demonstrating substantial performance improvement. Specifically, our image quality assessment method ranked first on the SAUD2.0 dataset and nearly achieved first place on the SUID2021 dataset, with only a slight margin separating it from the top performance. For the UWIQA dataset, our method ranks second. Overall, there is a noticeable trend of performance improvement as the dataset size increases, moving from UWIQA to SUID2021 and then to SAUD2.0. This trend in performance enhancement is closely related to the dataset size, which is attributable to the use of deep learning techniques in our proposed method. The effectiveness of deep learning models is typically influenced by the availability of large quantities of high-quality data.
@@ 4.3 Ablation Experiments
5 Conclusion
In this paper, we propose a novel physically imaging-guided framework for UIQA (PIGUIQA). By considering both direct transmission attenuation and backwards scattering effects, we integrate physical prior knowledge of the underwater imaging process into the deep learning-based model. The incorporation of advanced physics-based estimation allows for the definition of distortion metrics that capture the impact of these phenomena on visual perception. Moreover, we propose local perceptual modules based on the NA mechanism, which enhances the model’s ability to capture subtle image features and adaptively assess distortions in a region-specific manner. The experimental results demonstrate that the PIGUIQA not only outperforms existing methods in UIQA but also exhibits robust generalizability.
[…]
Supplementary Material
6 Visualization of Correlation Performance
@@
7 Visualization of Local Distortion-Aware Functions

### v2

# PIGUIQA: A Physical Imaging Guided Perceptual Framework for Underwater Image Quality Assessment

@@ Abstract
In this paper, we propose a Physical Imaging Guided perceptual framework for Underwater Image Quality Assessment (UIQA), termed PIGUIQA. First, we formulate UIQA as a comprehensive problem that considers the combined effects of direct transmission attenuation and backward scattering on image perception. By leveraging underwater radiative transfer theory, we systematically integrate physics-based imaging estimations to establish quantitative metrics for these distortions. Second, recognizing spatial variations in image content significance and human perceptual sensitivity to distortions, we design a module built upon a neighborhood attention mechanism for local perception of images. This module effectively captures subtle features in images, thereby enhancing the adaptive perception of distortions on the basis of local information. Third, by employing a global perceptual aggregator that further integrates holistic image scene with underwater distortion information, the proposed model accurately predicts image quality scores. Extensive experiments across multiple benchmarks demonstrate that PIGUIQA achieves state-of-the-art performance while maintaining robust cross-dataset generalizability. The implementation is publicly available at https:// anonymous.4open.science/r/PIGUIQA-A465/.
@@ 1 Introduction
The underwater world is rich in resources, and underwater images provide a crucial medium for exploring environments such as oceans and lakes by accurately and intuitively capturing underwater information [18]. However, the complexities of the underwater imaging environment often result in suboptimal image quality [34, 13, 48]. Therefore, underwater image quality assessment (UIQA) is fundamental for underwater image processing. UIQA methods help determine whether underwater images meet quality requirements for various applications and assess the visual quality of enhanced underwater images.
According to the accessibility of ideal references, image quality assessment (IQA) methods can be broadly categorized into full-reference (FR) [40], reduced-reference (RR) [49] and no-reference (NR) [30] approaches. For underwater images, obtaining perfect undistorted references is often impractical; thus, typical FR-IQA approaches, such as the structural similarity index measure (SSIM) [36] and feature similarity indexing method (FSIM) [44], are not applicable. Although NR-IQA approaches [23, 6, 26] have been developed for many years, most of them are designed to evaluate generic images. Many NR-IQA methods [19, 25, 17] are based on statistical properties, such as natural scene statistics (NSS), which are typically observed under good lighting conditions and in clear visual environments. However, the unique characteristics of underwater environments result in light propagation and scattering rules that are significantly different from those in terrestrial natural scenes. As a result, the statistical properties of NSS often fall short in accurately describing and assessing the quality of underwater images.
@@
Another primary challenge is that underwater images differ markedly from terrestrial images because of their specific imaging environments and lighting conditions. As a result, traditional NR-IQA methods that perform well in terrestrial image processing cannot be directly applied to underwater images. Specifically, as shown in Fig. 1, the underwater environment is complex and variable, with light propagation affected by scattering and absorption. Different wavelengths of light attenuate at different rates, with longer wavelengths diminishing more rapidly in water. Consequently, shorter wavelengths such as green and blue light can travel further underwater, causing underwater images to exhibit a blue–green hue and suffering from significant color distortion. The scattering of light in water can be categorized into forward scattering and backward scattering on the basis of the angle. Forward scattering occurs when light reflected from the target object deviates from its original path before it reaches the sensor, resulting in blurred underwater images. Backward scattering, on the other hand, involves a large amount of stray light entering the sensor, leading to visual issues such as loss of detail, low overall contrast, and increased background noise in underwater images. The need for the development of specialized UIQA methods that can effectively address the unique challenges posed by underwater imaging is urgent.
@@
To address the aforementioned issues, combining prior knowledge of underwater imaging with the excellent feature extraction capabilities of deep learning presents an effective solution. On this basis, we propose a physically imaging guided framework for UIQA (PIGUIQA) by integrating the physical model of underwater light propagation with deep learning techniques. Specifically, the framework considers distortions caused by backward scattering and direct transmission processes. These factors play crucial roles in underwater imaging, influencing image clarity, contrast, and color representation. Within the PIGUIQA framework, we leverage the powerful feature extraction capabilities of deep learning models to capture both local details and global perceptual features in images. Capturing local details allows the model to perceive distortions in critical regions of underwater images, such as object edges and texture information, whereas global perceptual features help the model understand the overall structure and context of the entire image. The main contributions of this work can be summarized as follows:
@@
  • We formulate UIQA as a comprehensive problem that accounts for the combined effects of direct transmission attenuation and backward scattering on image perception. We integrate advanced physics-based underwater imaging estimations into our framework. From the perspective of matrix theory, we define distortion metrics to measure these impacts.
@@
  • We design a neighborhood attention (NA)-based local perceptual module to capture subtle features in images, thereby enhancing the adaptive perception of distortions via local information. Additionally, a global perceptual aggregator is employed to integrate the original image content with underwater image distortion information, assisting the model in understanding the overall structure and context of the entire image.
@@ 3.1 Problem formulation
For underwater images, we specifically analyse image quality by considering the causes of distortion inherent to underwater imaging. In real underwater scenarios, the camera is often positioned close to the underwater scene. Consequently, the impact of forward scattering on image quality can be neglected, resulting in only direct transmission attenuation and backward scattering components to be considered. This relationship is represented as:
@@
where $Q$ is the ground-truth quality of an underwater image and where $d_{1}$ and $d_{2}$ represent the distance maps measuring transmission attenuation and backward scattering distortion between $T$ and $E$ and between $B$ and $O$, respectively. $f_{1}$ and $f_{2}$ are local distortion-aware functions to be learned, and $f_{3}$ is a global perceptual function to be learned. This formulation captures the essential aspects of UIQA by integrating a physical model into the evaluation process.
@@ 3.2 Framework
3.3 Underwater Imaging Distortion Measurements
A precise and detailed version of Eq. (2) is as follows [1]:
@@ 3.3 Underwater Imaging Distortion Measurements
where $c\in\{R,G,B\}$ is the color channel; $I_{per}$ and $I_{dis}$ are distortion-free and distorted images, respectively; $l$ is the object‒camera distance; $B_{\infty}$ is the background ambient light image; and $\beta_{D}$ and $\beta_{B}$ are the direct attenuation and backward scattering coefficients, respectively.
Based on Equation (2), many underwater image restoration methods [2, 38, 34] utilizing imaging estimation have been proposed, which can be mathematically formulated as:
$\hat{T},\hat{B},\hat{I}_{per}=\mathcal{F}(I_{dis})$ (5)
where $\mathcal{F}(\cdot)$ denotes the restoration model which can estimate the transmission attenuation matrix $T$ and background clutter matrix $B$, as defined in Eq. (2). Given that the primary objective of this study is to propose an UIQA framework rather than developing underwater image restoration algorithms which constitutes a distinct and complex research topic in computer vision, we employ a pre-trained network to handle the imaging estimation component.
Figure 3: Illustration of underwater imaging estimation via Eq. (5). (a): Input distorted underwater images $I_{dis}$. (b): Estimated transmission attenuation map $\hat{T}$. (c): Estimated background clutter map $\hat{B}$. (d): Restored underwater images $\hat{I}_{per}$. In the transmission attenuation map, darker areas signify more severe scattering and greater distortion, whereas in the background clutter map, brighter areas indicate a greater degree of background noise.
As demonstrated in Fig. 3, underwater imaging estimation enables quantitative characterization of distortion patterns, providing critical a priori information for subsequent image quality assessment. Considering the varying depths and optical properties of different regions and objects in underwater images, we decompose the image transmission attenuation map $T$ and the background clutter map $B$ into patches of size $N\times N$. This strategy facilitates the model’s learning of local features, thereby enhancing the perception of details in images with intricate content. Consequently, we address the design of imaging distortion metrics for each patch.
@@
The background clutter map can be regarded as an additional noise overlay on the original image, arising from background textures, variations in lighting, or the presence of nontarget objects. According to the theory of human visual sensitivity, the HVS can only distinctly recognize distortion when the intensity of the noise surpasses a specific perceptual threshold. Thus, we propose a backward scattering distortion measurement for the $k$-th patch of $B$:
@@ 4.1 Experimental Settings
WaDIQaM [3] 0.8565 0.8408 0.6617 10.8477 0.5099 0.4876 0.3414 1.8250
DEIQT [29] $\mathbf{0.9114}$ $\mathbf{0.9007}$ $\mathbf{0.7272}$ $\mathbf{8.5378}$ $\mathbf{0.5719}$ $\mathbf{0.5709}$ $\mathbf{0.4028}$ $\mathbf{1.8202}$
B-FEN [39] 0.8512 0.8441 0.6595 10.8642 0.4841 0.4766 0.3261 1.8564
@@
Implementation details. The proposed FIGUIQA network is trained ***The implementation code is programmed in PyTorch and is publicly available at https://anonymous.4open.science/r/PIGUIQA-A465/. The experiments are run on a PC with a Linux operating system and are configured with an NVIDIA GeForce RTX 4090 GPU with 24 GB of memory. using the default settings of the Adam optimizer. The learning rate is set to $10^{-4}$. To ensure the stability of the final model, a total of 1000 epochs are conducted during training. To prevent overfitting, data augmentation techniques such as random horizontal and vertical flipping, as well as rotation within the range of $[-15,15]$ degrees, are employed. The size of patches is set to $16\times 16$, which is a common setting in computer vision. SyreaNet [38] is selected as the pre-trained underwater imaging estimation model in the framework based on its state-of-the-art performance.
@@ 4.2 Performance Comparison
We conducted a systematic comparative analysis with 28 commonly-used or state-of-the-art no-reference image quality assessment (NR-IQA) and UIQA methods, including BRISQUE [23], BLIINDSII [31], NFERM [6], NIQE [24], IL-NIQE [45], SNP-NIQE [20], BIQME [7], FRIQUEE [5], PIQE [26], NRSL [17], RISE [16], DIIVINE [25], NPQI [19], dipIQ [22], HyperIQA [32], WaDIQaM [3], DEIQT [29], UNIQUE [47], DBCNN [46], B-FEN [39], UCIQE [41], UIQM [27], CCF [35], FDUM [43], NUIQ [14], MCOLE [15], Twice-Mix [4], and UIQI [21]. To ensure the reliability and stability of the proposed NAFRAD method, we repeat the training and testing process 10 times with randomized data partitioning, using the average of the test results as the final metric.
@@ 4.2.1 Single-Dataset Scenarios
In the experimental scenario involving single datasets, we randomly selected 80% of the data from each dataset for training and used the remaining 20% for evaluation. The experimental results are presented in Tables 1, 2, and 3. Our proposed method achieves the highest performance on the SAUD2.0 dataset, significantly surpassing the second-place method, MCOLE [15], demonstrating substantial performance improvement. Specifically, our image quality assessment method ranked first on the SAUD2.0 dataset and nearly achieved first place on the SUID2021 dataset, with only a slight margin separating it from the top performance. For the UWIQA dataset, our method ranks second. Overall, there is a noticeable trend of performance improvement as the dataset size increases, moving from UWIQA to SUID2021 and then to SAUD2.0. This trend in performance enhancement is closely related to the dataset size, which is attributable to the use of deep learning techniques in our proposed method. The effectiveness of deep learning models is typically influenced by the availability of large quantities of high-quality data.
@@ 4.3 Ablation Experiments
4.4 Visualization of Correlation Performance
@@
4.5 Visualization of Local Distortion-Aware Functions
@@
5 Conclusion
In this paper, we propose a novel physical imaging guided framework for UIQA, PIGUIQA. By modeling underwater image degradation as a combined effect of direct transmission attenuation and backward scattering, we establish physics-based distortion metrics that quantitatively characterize their perceptual impacts. Moreover, local perceptual modules utilizing a neighborhood attention mechanism are designed to capture subtle spatial features and prioritize perceptually sensitive areas. The framework also introduces a global perceptual aggregator that holistically fuses scene semantics with distortion characteristics to ensure accurate quality score prediction. Extensive experiments across diverse underwater benchmarks validate that PIGUIQA achieves state-of-the-art performance, demonstrating superior correlation with human subjective judgments and exceptional cross-dataset generalization capabilities.